# Network Analytics for Anti-Money Laundering – A Systematic Literature Review and Experimental Evaluation

### v2

@@ 4.1 Model Specification
To set up the experimental evaluation, we start from the wider literature on network representation. Extensive overviews are given in [20, 52, 47], classifying the methods according to their input, output, and embedding technique. A sub-branch of representation learning applies deep learning methods [141, 127]. Based on these papers, we select the methods that are most prominently used. The data set has labels at node level, our experiments will be set up as supervised node classification. We refined the selection by considering the fact that, for AML, transaction data contains networks with millions of nodes and edges, making scalability an important selection criterion. Table 7 provides an overview of the methods selected. They are grouped into three categories [20, 52, 47]: manual feature engineering, shallow representation learning and deep representation learning. These methods will be tested against a baseline model that only includes the intrinsic features, denoted by IF.
@@
Manual feature engineering Density [12]: Number of edges in the egonet relative to the maximal number possible.
@@
PageRank [95]: Score indicating the importance of a node.
Shallow representation learning DeepWalk [96]: Random walk through the network. Embedding via word2vec.
Node2vec [48]: Truncated random walks (breadth-first vs. depth-first). Embedding via word2vec.
Deep representation learning GCN [65]: Graph convolutional network
GraphSAGE [52]: Aggregation based on fixed-size sample of neighbours
GAT [120]: Graph attention network
GIN [132]: Graph isomorphism network
Table 7: Network Analytics Methods. The methods used to generate network features for the experimental evaluation.
Manual feature engineering is the first category of methods and includes both local and global metrics. The first is the density of the ego-network. A node’s ego-network consists of said node, its direct neighbours and all connections between the neighbours. The density is the relative number of connection in this ego-network compared to the theoretical maximum. Summary statistics, i.e., minimum, mean and maximum values, of the density of the node’s neighbours are also included.
A second type of manual features are centrality metrics, quantifying the global position of the node in the network [90]. The most popular centrality metrics are the betweenness, closeness and eigenvector centrality. The betweenness quantifies the number of shortest paths between any two nodes in the network passing through a given node. The closeness measures how close a node is to all other nodes, and is the inverse of the average distance to the other nodes. The eigenvector centrality is high for nodes that are connected (but are not necessarily themselves) to important nodes in the network. The importance of the node itself can be quantified using the PageRank [95].
@@
Node2vec [48] extends the capabilities of the random walks by introducing two additional hyperparameters, $p$ and $q$, resulting in the option to either use breadth-first or depth-first sampling. The method samples each node according to the following unnormalized transition probability, $\alpha$, at node $v$:
@@
Deep network representation learning adapts deep learning methods to be directly applicable to networks. This paper applies different graph neural network (GNN) architectures. The first graph neural network framework was introduced in [103, 102] and is based on recurrent neural networks. The idea is that the network defines how nodes can propagate information to each other. This propagation step is iterated until a state of convergence. As noted in [52], the initial message passing idea is mostly used for sub-network embedding calculation.
@@
Neighbour aggregation has seen multiple extensions, two of which this paper will evaluate. The first is GraphSAGE [52], which samples a fixed number of neighbours for each node, to make the calculations scalable to large graphs. First, the representations of the sampled neighbours are aggregated, with the authors proposing the mean, LSTM and pooling aggregators [52]. Then, the aggregated neighbourhood representation is concatenated with the node’s representation. Finally, the representation is updated using the weights and activation function. For a given node $v$, this is calculated as:
@@
The second extension is graph attention network (GAT) [120], which adds attention mechanism when learning the representations. Our work uses the modified version of the extension mechanism, introduced to fix the static attention problem [18]:
@@
The above formulas apply one attention mechanism. As mentioned in [120], having multiple mechanisms - also called heads - at the same step can improve stability. We will follow the method in [120], where the intermediate steps concatenate the results over the heads, while the final layer averages the results.
@@
Figure 3: Illustration feature engineering. For the feature engineering and shallow representation methods, the obtained network features/embedding is added to the intrinsic ones.
Different decoders are used for the final classification task. A deep neural network with two hidden layers, each of dimension ten is used for the manual features and shallow representation methods. For the graph neural networks, the GNN layers are followed by a single linear layer for making predictions.
@@ 4.2 Data
We use the Elliptic data set [1, 122], which contains Bitcoin transactions for 49 time steps each being around two weeks long. The network has 203 769 nodes and 234 355 edges. The nodes correspond to transactions, and an edge represents that the output of one transaction is the input of the next. Additionally, the data set contains 166 pre-calculated features. These are split into 94 transaction-specific, or local, features and 72 aggregated features, summarizing the local features of a node’s neighbours.
The Elliptic data set has only numerical features, which have already been standardised. The local features will be used for all methods. However, since the deep representation learning methods aggregate the neighbourhood information, including the node features, we do not include the aggregated features for these. This was also addressed by the authors [122], stating that graph neural networks are better to address the heterogeneity of the neighbourhoods than the aggregated features.
Around $33\%$ of transactions are labelled, of which 4 545 are illicit and 42 019 are licit. Hence, the classification problem is imbalanced with around $11\%$ of labelled nodes, or $2\%$ of all nodes labelled illicit. We note that an illicit label does not automatically mean that the transaction is used to launder money. However, criminals will try to obscure the source of these illicit funds, so we believe that it is still informative.
@@ 4.3 Hyperparameter Tuning
Most methods are highly dependent on their hyperparameters. Hence, finding those that give the best performance is crucial. However, it is not feasible to tune these using a grid search. As the methods often have multiple hyperparameters, the number of combinations increases exponentially. Transaction networks are very large, which results in a relative long training time for each individual combination of hyperparameters. Therefore, we tune the hyperparameters efficiently via Optuna [5], whereby we give a range of possible values to select from. We relied on optuna’s default TPESampler sampler. After these trials, the combination is retained that maximises the AUC-PR on the validation set. The ranges are summarised in Table 8. These are based on the ones often used in the literature. A model only based on the intrinsic features (IF) is included. The number of trials is set to 50 for DeepWalk and node2vec, and 100 for the GNNs.
Hyperparameter Tuning Range Model (hyperparameter value)
$\alpha$: random jump parameter $[0.1,0.9]$ PageRank (0.593)
Number of walks per node $[1,3]\cap\mathbb{N}$ DeepWalk (2), node2vec (1)
Walk length $[3,10]\cap\mathbb{N}$ Deepwalk (3), node2vec (9)
Word2vec context window size $[2,10]\cap\mathbb{N}$ DeepWalk (2), node2vec (5)
Latent dimension $[2,64]\cap\mathbb{N}$ DeepWalk (5), node2vec (47)
Latent dimension $[32,128]\cap\mathbb{N}$ GCN (87), GraphSAGE (77), GAT (104), GIN (98)
$p$: return parameter $[0.5,2]$ node2vec (1.17)
$q$: in-out parameter $[0.5,2]$ node2vec (1.60)
Number of negative samples $[1,5]\cap\mathbb{N}$ DeepWalk (1), node2vec (1)
GNN hidden dimensions $[64,256]\cap\mathbb{N}$ GCN (217), GraphSAGE (192), GAT (N/A), GIN (N/A)
GNN layers $[1,3]\cap\mathbb{N}$ GCN (3), GraphSAGE (1), GAT (1), GIN (1)
Number of neighbourhood samples $[2,5]\cap\mathbb{N}$ GraphSAGE (2)
Learning rate $[0.01,0.1]$ IF (0.0163), Manual (0.0166), DeepWalk (0.0554), node2vec (0.0159), GCN (0.0864), GraphSAGE (0.0690), GAT (0.0487), GIN (0.0452)
Aggregator {min, mean, max} GraphSAGE (max)
Number of heads $[1,5]\cap\mathbb{N}$ GAT (1)
Dropout rate $[0,0.5]$ GCN (0.057), GraphSAGE (0.345), GAT (0.471), GIN (0.384)
Number of layers decoder $[1,3]\cap\mathbb{N}$ IF (1), Manual (1)
Hidden dimension decoder $[5,20]\cap\mathbb{N}$ IF (5), Manual (6)
Number of epochs decoder $[5,500]\cap\mathbb{N}$ IF (497)
Number of epochs decoder $[5,100]\cap\mathbb{N}$ Manual (64), DeepWalk (80), node2vec (93)
Number of epochs $[5,500]\cap\mathbb{N}$ DeepWalk (176), node2vec (222), GCN (174), GraphSAGE (494), GAT (282), GIN (42)
@@
Table 8: Hyperparameter tuning. The overview of the hyperparameter values. The tuned hyperparameter value is added between parentheses next to the name of the model.
@@
The decoders used for classification all have a fixed architecture (Section 4.1). For the GNNs, the training of the decoder weights happens simultaneously as the other weights during training. For the manual and shallow representation features, the decoder/classifier needs to be trained afterwards. The only hyperparameter tuned is the number of epochs. The results are presented in parentheses in Table 8.
@@ 4.4 Performance Metrics
We tackle money laundering as a binary classification problem. We adopt the most popular performance metrics from the literature, namely the precision, recall and F1-score (Section 3.1). A threshold is needed for the output of the model to classify an instance as a money launderer. Since investigation resources are limited [118], the thresholds will typically be set relatively high, e.g., by classifying the top $0.1\%$, $1\%$ or $10\%$ of scores as money laundering. This is supplemented by a threshold equal to the relative prevalence of money laundering in the data set. Here, we take $11\%$, since we are only considering labelled nodes in the training and testing.
@@
These metrics are not calculated once, but multiple times on the test set to get a grip on the variance of the results. This variance is determined by inserting randomness in the observations in the test set as follows. The manual network features and shallow representation embeddings result in classic tabular features (cf. Figure 3). For these, we apply a bootstrap method, where a new test set is constructed of the same size as the original, but where the observations are sampled from the original test set with replacement. This is slightly different for the deep representation methods. In the implementation, test masks are used. Therefore, we initialise new test masks by randomly selecting half the test masks from the original ones.
@@ 5 Results and Discussion
The results for the threshold-independent metrics, i.e., AUC-ROC and AUC-PR, on the test set are shown in Table 9, including the standard deviation calculated as described in Section 4.4. Additionally, the threshold-dependent results, i.e., precision, recall and F1-score, are given Table 10 for the top-$1\%$ results. To analyse the sensitivity to the threshold, the metrics are calculated when classifying the top $0.1\%$, $10\%$ and $p\%$ of scores as money laundering, with $p\%$ indicating the relative prevalence of the money laundering label in the training data set. The results for these three thresholds are included in E (in Figure 12 and 3, respectively).
Elliptic
Methods AUC-ROC AUC-PR
Intrinsic features $0.7632\pm 0.0131$ $0.5628\pm 0.0216$
Egonet features $0.8217\pm 0.0100$ $0.5417\pm 0.0210$
Deepwalk-NI $0.4500\pm 0.0114$ $0.0488\pm 0.0014$
Deepwalk $0.7425\pm 0.0122$ $0.5474\pm 0.0192$
Node2vec-NI $0.5263\pm 0.0108$ $0.0594\pm 0.0024$
Node2vec $0.7600\pm 0.0125$ $0.5553\pm 0.0188$
GCN $\boldsymbol{0.8329\pm 0.0099}$ $0.5946\pm 0.0184$
GraphSAGE $0.8279\pm 0.0005$ $\boldsymbol{0.6312\pm 0.0009}$
GAT $0.8102\pm 0.0111$ $0.5886\pm 0.0193$
GIN $0.8089\pm 0.0111$ $0.5517\pm 0.0212$
Table 9: Threshold-independent metrics: AUC-ROC and AUC-PR values over the different methods for the Elliptic data set, based on the test set. The standard deviation is also reported.
Elliptic

### v3

@@ 4.1 Model Specification
To set up the experimental evaluation, we start from the wider literature on network representation. Extensive overviews are given by Cai et al. [18], Hamilton et al. [55] and Goyal and Ferrara [50], classifying the methods according to their input, output, and embedding technique. A sub-branch of representation learning applies deep learning methods [158, 144]. Based on these papers, we select the methods that are most prominently used. We refined the selection by considering that, for AML, transaction data contains networks with millions of nodes and edges, making scalability an important selection criterion. Table 8 provides an overview of the methods selected. They are grouped into three categories [18, 55, 50]: manual feature engineering, shallow representation learning and deep representation learning. These methods will be tested against a baseline model that only includes the intrinsic features, denoted by IF.
Table 8: The methods used to generate network features for the experimental evaluation.
@@
Manual feature engineering Density Baesens et al. [10]: Number of edges in the egonet relative to the maximal number possible.
@@
PageRank Page et al. [104]: Score indicating the importance of a node.
Shallow representation learning DeepWalk Perozzi et al. [105]: Random walk through the network. Embedding via word2vec.
Node2vec Grover and Leskovec [51]: Truncated random walks (breadth-first vs. depth-first). Embedding via word2vec.
Deep representation learning GCN Kipf and Welling [71]: Graph convolutional network
GraphSAGE Hamilton et al. [55]: Aggregation based on fixed-size sample of neighbours
GAT Veličković et al. [137]: Graph attention network
GIN Xu et al. [149]: Graph isomorphism network
Manual feature engineering includes both local and global metrics. The first is the density of the ego-network. A node’s ego-network consists of said node, its direct neighbours and all connections between the neighbours. The density is the relative number of connections in this ego-network compared to the theoretical maximum. Summary statistics, i.e., minimum, mean and maximum values, of the density of the node’s neighbours are also included.
A second type of manual features are centrality metrics, quantifying the global position of the node in the network [99]. The most popular centrality metrics are the betweenness, closeness and eigenvector centrality. The betweenness quantifies the number of shortest paths between any two nodes in the network passing through a given node. The closeness measures how close a node is to all other nodes, and is the inverse of the average distance to the other nodes. The eigenvector centrality is high for nodes that are connected to (but are not necessarily themselves) important nodes in the network. The importance of the node itself can be quantified using the PageRank [104].
@@
Node2vec [51] extends the capabilities of the random walks by introducing two additional hyperparameters, $p$ and $q$, resulting in the option to either use breadth-first or depth-first sampling. The method samples each node according to the following unnormalized transition probability, $\alpha$, at node $v$, when coming from node $t$:
@@
The main drawbacks of shallow representation learning methods, are that (1) they are transductive, so need to be retrained fully when a new node is added to the network, and (2) retraining on the same network often results in a completely different embedding. This second point can be understood intuitively by viewing two embeddings of the same network, but where one is a rotation of the other. Both are equally good, but a downstream classifier trained on one cannot be used to make predictions on the other, as the coordinates, i.e., the features values, are completely different.
This needs to be considered when training and comparing models. In the experiments presented below, we construct the embedding first on the network with nodes in the training and validation set, only training on the labels from the training set. Afterwards, the embedding is constructed for the full network, and we train a new classifier using the labels of the train and validation set.
This mimics the applications of these methods in reality, since a bank will have the full transaction network up to the current time, with some historical labels from past moments. Based on this, they need to determine which accounts to investigate next for money laundering. This also highlights the difference with credit card fraud, where a new transaction comes in and a decision needs to be made in only a few seconds, limiting the ability to retrain the embedding on the whole network.
Deep network representation learning adapts deep learning methods to be directly applicable to networks. This paper applies different graph neural network (GNN) architectures. The first graph neural network framework was introduced by Scarselli et al. [113] and Scarselli et al. [114] and is based on recurrent neural networks. The idea is that the network defines how nodes can propagate information to each other. This propagation step is iterated until a state of convergence. As noted by Hamilton et al. [55], the initial message passing idea is mostly used for sub-network embedding calculation.
@@
Neighbour aggregation has seen multiple extensions, two of which this paper will evaluate. The first is GraphSAGE [55], which samples a fixed number of neighbours for each node, to make the calculations scalable to large graphs. The representations of the sampled neighbours are aggregated, with the authors proposing the mean, LSTM and pooling aggregators [55]. Then, the aggregated neighbourhood representation is concatenated with the node’s representation. Finally, the representation is updated using the weights and activation function. For a given node $v$, this is calculated as:
@@
The second extension is graph attention network (GAT) [137], which adds an attention mechanism when learning the representations. Our work uses the modified version of the attention mechanism, introduced to fix the static attention problem [16]:
@@
The above formulas apply one attention mechanism. As mentioned by Veličković et al. [137], having multiple mechanisms - also called heads - can improve stability. We will follow the method of Veličković et al. [137], where the intermediate steps concatenate the results over the heads, while the final layer averages the results.
@@
Figure 3: For the feature engineering and shallow representation methods, the obtained network features/embedding is added to the intrinsic ones.
Different decoders are used for the final classification task in the supervised setting. A deep neural network with two hidden layers, each of dimension ten is used for the manual features and shallow representation methods. For the graph neural networks, the GNN layers are followed by a single linear layer for making predictions.
For the unsupervised model, we apply isolation forest [82], as it achieves state-of-the-art performance for outlier/anomaly detection [126, 65]. This will be applied to the intrinsic and manual features, and the shallow representations, since the isolation forest requires a tabular input. As mentioned before, a larger experiment comparing unsupervised methods for network analytics is left for future work.
@@ 4.2 Data
For the evaluation of the different methods, we rely on two data sets; the elliptic [37, 139] and IBM money laundering [6, 36] data sets. Using these two, we cover the two main lines of research in the literature. The elliptic data set concerns real-world crypto-transactions, while the IBM data set deals with simulated transactions in fiat currencies. We also discuss the main characteristics of and differences between the two networks based on different metrics.
4.2.1 Elliptic.
The Elliptic data set [37, 139] contains Bitcoin transactions for 49 time steps each being around two weeks long. We make this choice since (1) we know from Section 3.2 that this data set is widely used and (2) the instances have a ground truth label, facilitating evaluation of the methods. The network has 203 769 nodes and 234 355 edges. The nodes correspond to transactions, and an edge represents that the output of one transaction is the input of the next. Additionally, the data set contains 166 pre-calculated features. These are split into 94 transaction-specific, or local, features and 72 aggregated features, summarizing the local features of a node’s neighbours.
The Elliptic data set has only numerical features, which have already been standardised. The local features will be used for all methods. However, since the deep representation learning methods aggregate the neighbourhood information, including the node features, we do not include the aggregated features for these. This was also addressed by Weber et al. [139], stating that graph neural networks are better to address the heterogeneity of the neighbourhoods than the aggregated features.
Around $33\%$ of transactions are labelled, of which 4 545 are illicit and 42 019 are licit. Hence, the classification problem is imbalanced with around $11\%$ of nodes labelled, or $2\%$ of all nodes labelled illicit. We note that an illicit label does not automatically mean that the transaction is used to launder money. However, criminals will try to obscure the source of these illicit funds, so we believe that it is still informative.
@@
4.2.2 IBM Money Laundering.
The second data set is the HI-Small data set, taken from Kaggle, consisting of simulated transactions in a virtual world between individuals, companies and banks [6, 36]. A fraction of these transactions are labelled as money laundering. Given that the labels are at transaction level, the network is constructed with the transactions as nodes. Edges are added from transaction $v_{i}$ to $v_{j}$ if the receiver of transaction $v_{i}$ is the sender of transaction $v_{j}$, transaction $v_{j}$ happened before $v_{i}$, and the difference in time is smaller than $\Delta t$, i.e., $\Delta t\geq t_{j}-t_{i}\geq 0$. This is inspired by the work by Tariq and Hassani [124], and allows to capture the money flow. We take a 60-20-20 train-validation-test split based on transaction time.
Since this results in a massive data set and tuning all hyperparameter for DeepWalk and node2vec would take too long, we only take the last 500 000 transactions, and set $\Delta t$ to four hours. This results in a network with 500 000 nodes and 1 278 952 edges. Just as with the elliptic data set, this subsample of the IBM money laundering data set is highly imbalanced. Just 1356 transactions, i.e., $0.27\%$, are labelled as money laundering.
Each transaction has a couple of features, of which we use the following: ‘Amount Received’, ‘Receiving Currency’, ‘Amount Paid’, ‘Payment Currency’, ‘Payment Format’. Next to that, we extract the day of the week, hour and minute out of a transaction’s date stamp. The categorical features, i.e., ‘Receiving Currency’, ‘Payment Currency’, ‘Payment Format’, are transformed to numerical using one-hot-encoding. In the remainder of this paper, we denote this data set as IBM-AML.
An analysis of the topology of both networks is given in Table 9, based on well-known network features. IBM-AML seems to be much more connected, since the average degree is higher, and the average path length is lower compared to the Elliptic data set. This is corroborated by the degree distribution and the distributions of the closeness centrality, in Figure 4 and Figure 5, respectively. We see that the Elliptic data set more closely resembles a scale-free distribution, while the IBM-AML data set deviates from this with fewer nodes with low degree and more nodes with higher degree (hubs). The existence of hubs resulted in a much lower average path length, and higher closeness centrality values.
Table 9: Network features for the Elliptic and IBM-AML data set.
Network Avg. Degree Avg. Path Length Clustering
Elliptic 2.3 15.96 0.0117
IBM-AML 5.1 3.78 0.0000
Figure 4: Comparison of the degree distributions of the Elliptic and IBM-AML data sets.
Figure 5: Comparison of the distributions of the closeness centrality of the Elliptic and IBM-AML data sets.
@@ 4.3 Hyperparameter Tuning
Most methods are highly dependent on their hyperparameters. Hence, finding those that give the best performance is crucial. However, it is not feasible to tune these using a grid search. As the methods often have many hyperparameters, the number of combinations increases exponentially. Additionally, transaction networks are very large, which results in a relative long training time for each individual combination of hyperparameters. Therefore, we tune the hyperparameters efficiently via Optuna [2], whereby we give a range of possible values to select from. We use optuna’s default TPESampler sampler.
After these trials, the combination is retained that maximises the AUC-PR on the validation set. The ranges are summarised in Table 10 and 11, for supervised and unsupervised learning, respectively. These are based on the ones often used in the literature. As the embeddings from the shallow representation methods are used in both a supervised and unsupervised setting, the hyperparameters will be tuned for the two applications separately. This will be indicated with superscript S (for supervised) and U (for unsupervised) when presenting the results. A model only based on the intrinsic features (IF) is included. The number of trials is set to 50 for DeepWalk and node2vec, and 100 for the GNNs.
Since the IBM data set is much larger, scalability in both time and memory are becoming very important. Especially DeepWalk and node2vec take quite some time to run. The higher average degree also makes transition probability calculations more expensive. That is why the number of epochs for calculating the embedding are capped at 100. The attention mechanism of GAT is conditional on the query node, meaning that attention is learned for each node pair. To be able to train the model in memory, we limit the attention head to a maximum of two.
@@
Table 10: The overview of the hyperparameter values for the supervised learning models. The tuned hyperparameter value is added between parentheses next to the name of the model.
Hyperparameter Tuning Range Model (hyperparameter value Elliptic $\mid$ IBM )
[…]
Latent/Embedding dimension $[32,128]\cap\mathbb{N}$ GCN ($108\mid 97$), GraphSAGE ($48\mid 53$), GAT ($99\mid 51$), GIN ($74\mid 47$)
[…]
GNN layers $[1,3]\cap\mathbb{N}$ GCN ($1\mid 2$), GraphSAGE ($1\mid 2$), GAT ($1\mid 1$), GIN ($1\mid 1$)
Number of neighbourhood samples Not included at the moment. $[2,5]\cap\mathbb{N}$ GraphSAGE (2)
[…]
Number of attention heads $[1,5]\cap\mathbb{N}\mid[1,2]\cap\mathbb{N}$ GAT ($4\mid 2$)
Dropout rate $[0,0.5]$ GCN ($0.288\mid 0.163$), GraphSAGE ($0.350\mid 0.244$), GAT ($0.186\mid 0.188$), GIN ($0.271\mid 0.273$)
[…]
Table 11: The overview of the hyperparameter values for the unsupervised learning models. The tuned hyperparameter value is added between parentheses next to the name of the model.
Hyperparameter Tuning Range Model (hyperparameter value Elliptic $\mid$ IBM )
[…]
@@
The decoders used for classification all have a fixed architecture (Section 4.1). For the GNNs, the training of the decoder weights happens simultaneously as the other weights during training. For the manual and shallow representation features, the decoder/classifier needs to be trained afterwards. The only hyperparameter tuned is the number of epochs. The results are presented in parentheses in Table 10 and Table 11.
@@ 4.4 Performance Metrics
We tackle money laundering as a binary classification problem. We adopt the most popular performance metrics from the literature, namely the precision, recall and F1-score (Section 3.2). A threshold is needed for the output of the model to classify an instance as a money launderer. Since investigation resources are limited [133], the thresholds will typically be set relatively high, e.g., by classifying the top $0.1\%$, $1\%$ or $10\%$ of scores as money laundering. This is supplemented by a threshold equal to the relative prevalence of money laundering in the data sets.
@@
The performance metrics are calculated 100 times on a randomised version of the test set to get a grip on the variance of the results. The randomisation of the test set is done as follows. The manual network features and shallow representation embeddings result in classic tabular features (cf. Figure 3). For these, we apply a bootstrap method, where a new test set is constructed of the same size as the original, but where the observations are sampled from the original test set with replacement. The deep representation methods apply test masks. Therefore, we initialise new test masks by randomly selecting half the test masks from the original ones. Both ways of working result in 100 different test sets on which the performance metrics are calculated.
@@ 5 Results and Discussion
5.1 Supervised Learning
The results for the threshold-independent metrics, i.e., AUC-ROC and AUC-PR, on the test set are shown in Table 12, including the standard deviation calculated as described in Section 4.4. Additionally, the threshold-dependent results, i.e., precision, recall and F1-score, are given Table 13 and Table 14 for the top-$1\%$ results. To analyse the sensitivity to the threshold, the metrics are calculated when classifying the top $0.1\%$, $10\%$ and $p\%$ of scores as money laundering, with $p\%$ indicating the relative prevalence of the money laundering label in the training data set. The results for these three thresholds are included in Appendix E.
Table 12: Threshold-independent metrics: AUC-ROC and AUC-PR values over the different methods for the two data sets, based on the test set. The standard deviation is also reported.
Supervised Elliptic IBM-AML
Methods AUC-ROC AUC-PR AUC-ROC AUC-PR
Intrinsic features $0.877\pm 0.008$ $0.587\pm 0.022$ $0.755\pm 0.005$ $0.019\pm 0.001$
Egonet features $0.877\pm 0.008$ $0.604\pm 0.021$ $0.753\pm 0.006$ $0.018\pm 0.001$
DeepWalk $0.833\pm 0.009$ $0.562\pm 0.022$ $0.758\pm 0.005$ $\boldsymbol{0.019\pm 0.001}$
Node2vec $0.838\pm 0.010$ $0.622\pm 0.020$ $\boldsymbol{0.759\pm 0.005}$ $\boldsymbol{0.019\pm 0.001}$
GCN $0.862\pm 0.009$ $0.615\pm 0.020$ $0.572\pm 0.002$ $0.012\pm 0.000$
GraphSAGE $0.865\pm 0.008$ $0.637\pm 0.020$ $0.574\pm 0.002$ $0.011\pm 0.000$
GAT $\boldsymbol{0.892\pm 0.007}$ $\boldsymbol{0.679\pm 0.017}$ $0.267\pm 0.007$ $0.008\pm 0.000$
GIN $0.833\pm 0.010$ $0.551\pm 0.019$ $0.745\pm 0.006$ $0.018\pm 0.001$
Elliptic Supervised